OpenStudy (anonymous):

I have a pretty simple math problem: the price of a liter cola increased from $1.09 to $1.29. What was the percent of increase? I know to first subtract (1.29-1.09) that equals $.20. Next divide .20 by 1.09 (first move decimal points over) I got the same ander as the book, except it says the final answer is 18.3% increase. Where did the decimal come from? It's supposed to be in front of the 1. I think they're rounding, but how did the get that? And how are you supposed to round a non terminating, non repeating decimal(in finance)?

OpenStudy (kropot72):

The percentage increase is given by \[\large \frac{0.20}{1.09}\times\frac{100}{1}=18.3\ percent\] The answer is rounded to the nearest one-tenth of a percent.

OpenStudy (kropot72):

@Learea Do you find this helps your understanding?

OpenStudy (anonymous):

Thanks! Is that how numbers (with percents or money) is usually rounded? To the nearest one tenth?

OpenStudy (kropot72):

Well, the answer in the book has been rounded to the nearest one-tenth of a percent. However interest rates on investments are often stated to one-hundredth of a percent. Sometimes questions state what rounding should be given to an answer.

OpenStudy (anonymous):

Alright, thank you

OpenStudy (kropot72):

You're welcome :)